OpenStudy (anonymous):

Find the horizontal asymptote f(x)=(4t)/(3t^(2)+3)

OpenStudy (amistre64):

as t get very large; the bottom overcomes the top. such that we get a very large number on the bottom and a very small number up top; relatively speaking that is

OpenStudy (blockcolder):

\[\Large f(x)=\frac{4t}{3t^2+3}=\frac{\frac{4}{t}}{3+\frac{3}{t^2}}\] Now take the limit as t approaches infinity.

OpenStudy (amistre64):

to put it in persepctive; if you had 1/1000000000000 piece if a pie; youd have practically none

OpenStudy (amistre64):

blocks eq is fine ...

OpenStudy (anonymous):